6.042/18.062J Mathematics for Computer Science February 1, 2005 Srini Devadas and Eric Lehman

Course Information

Welcome to 6.042! In this course, we’ll teach you some mathematics that we think you’ll find useful in your study of computer science. This handout contains basic information about the class, but all this and more is also available on the course website:

         http://theory.csail.mit.edu/classes/6.042

Prerequisites. The only prerequisite is 18.01. If you have already taken 18.310 or 6.046, then you should not take 6.042.

Lecture. There are 90-minute lectures on Tuesday and Thursday in 32-123 at 11 AM.

Recitation. There are 1-hour recitations on Wednesday and Friday focused on solving problems in small groups. Attendance is required! The first recitation is tomorrow (Wednesday). We’ll post recitation assignments on the website tonight.

Office hours. Times and locations are on the website.

Reading. We’ll post lecture notes on the website after each class. The notes and lectures will be similar, but not identical. You are responsible for the content of both.

Homework. Typically, a problem set is released on the website on Tuesday, is due on the following Monday night at 9PM in 32-044, and is returned in recitation on Wednesday. Late homework is not accepted, but talk to your recitation instructor if a special situation arises.

Exams. There are three exams: two 2-hour quizes and a 3-hour final during finals week.

Collaboration. You are welcome to work with other students on homework, but your writeup must be entirely your own. Please do not refer to course materials from previous terms. On the top of your homework, list:

• all collaborators, other than course staff

• all written sources that you consulted aside from this term’s course materials

If you had no collaborators and consulted no written sources, then write, “I worked alone.” Homework without a collaboration statement will not be graded. Collaboration on exams is not allowed. If you somehow violate the collaboration policy, your best option is to tell us before we notice. Mistakes you confess are forgivable.

Grading We compute a percentage score based on your coursework and then assign a letter grade as follows:

A 88.0 - 100%
B 78.0 - 87.9%
C 65.0 - 77.9%
D 50.0 - 64.9%
F below 50%

Your percentage score is the weighted average of your scores in five areas: homework, recitation, quiz 1, quiz 2, and the final exam. Scores in the five individual areas are determined as follows:

Homework (25%) We drop your lowest score. We may normalize an entire recitation section upward, if necessary, to compensate for variations in grading standards.

Recitation (20%) Each recitation is worth 0, 1, or 2 points. If you attend for the full period and work constructively with your team, then you get 2 points. If you miss part of recitation or glaringly fail to work constructively with your team, then you get 1 point. If you are absent, you get 0 points. We drop your two lowest recitation scores.

Quiz 1 (20%), Quiz 2 (20%), Final (25%) We’ll cut 10% off the weight of your weakest exam.

Despite our best efforts, we sometimes make errors in grading. If you think your work on homework or exam was misgraded, please let us know. However, requests for regrades must be submitted within a week after the return date and before the final exam.

We reserve the right to assign a better course grade than the one determined by your scores.

How to Succeed in 6.042

We want everyone to get the best possible grade within the bounds of fairness. There is no curve; in principle, everyone could get an A. Here are some suggestions on how you can do well:

• Attend recitation! If you show up on time, stay the whole hour, and work constructively with teammates, then you’ve got 20% in the bag.

• Attend lecture! We’ll go over the major points of every topic and answer your questions. Lecture may cover examples not addressed in the notes.

• Collaborate with other students on the homework. Some problems in 6.042 are tricky and sharing insights can save you a lot of time.

• Rely on your recitation instructor. Attend office hours and send questions via email. (Office hours are also a good opportunity to find collaborators.)

• Exams are typically tough. The best preparation is to do your best on each homework and go over your mistakes afterward with your recitation instructor.

• If you are having one of those terms where you’re getting buried by 6.042, MIT, and life at large, come talk to us and we’ll see if we can help you out.

Good luck and we hope you enjoy the class!